# Amortized Bayesian Workflow

### v2

@@ 1 Introduction
In many statistical modeling applications, from finance to biology and neuroscience, we often aim to infer unknown parameters $\theta$ from observables $y$ modeled as a joint distribution $p(\theta,y)$ [e.g., 66, 76, 28, 40, 10, 49, 74, 86]. The posterior $p(\theta\,|\,y)$ is the statistically optimal solution to this inverse problem, and there are different computational approaches to approximate this target distribution.
Markov chain Monte Carlo (MCMC) methods constitute the most popular family of posterior sampling algorithms and still remain the gold standard for modern Bayesian inference due to their theoretical guarantees and powerful diagnostics [27, 26]. MCMC methods yield autocorrelated draws conditional on a fixed dataset $y_{\text{obs}}$. As a consequence, the probabilistic model has to be re-fit for each new dataset, which involves repeating the entire MCMC procedure from scratch. Modern implementations equip MCMC with state-of-the-art extensions, for example, through Hamiltonian dynamics (HMC; [56]), by minimizing the required tuning by users (NUTS; [37]), or by parallelizing thousands of chains on GPU hardware (ChEES-HMC; [36]). The well-established Bayesian workflow [27] leverages these tools in an iterative process of model specification, fitting, evaluation, and revision. While powerful, this approach becomes computationally burdensome when applied independently to large collections of datasets.
Differently, amortized Bayesian inference (ABI) aims to learn a direct mapping from observables $y$ to the corresponding posterior $p(\theta\,|\,y)$, using flexible function approximators such as deep neural networks [12, 62, 32, 60, 95, 77, 99]. Amortized inference typically follows a two-stage approach: (i) a training stage, where neural networks learn to distill information from the probabilistic model based on simulated examples of observations and parameters $(\theta,y)\sim p(\theta)\,p(y\,|\,\theta)$; and (ii) an inference stage where the neural networks approximate the posterior distribution for an unseen dataset $y_{\text{obs}}$ in near-instant time without repeating the training stage. In other words: The upfront training cost is amortized by negligible inference cost on arbitrary amounts of unseen test data. Owing to its reliance on simulated data, amortized inference in this form is a subset of simulation-based inference [12].
@@
However, amortized inference lacks the powerful diagnostics and gold-standard guarantees associated with MCMC samplers in the standard Bayesian workflow [27]. Yet, applying a standard workflow is computationally prohibitive at scale. In modern Bayesian computation, MCMC and ABI occupy different ends of a Pareto frontier (see Figure 1): the former provides reliable accuracy at high cost, while the latter offers near-instant inference speed with limited per-dataset reliability [34, 71, 48].
In this paper, we propose an adaptive workflow that consistently yields high-quality posterior draws while remaining computationally efficient. Our proposed workflow moves along the Pareto front, enabling fast-and-accurate inference when possible, and slow-but-guaranteed-accurate inference when necessary (see Figure 1). It combines the strengths of ABI and MCMC by incorporating diagnostic checks to guide inference decisions and reuse computations wherever possible. The resulting amortized Bayesian workflow therefore offers a principled, scalable, and diagnostic-driven approach for efficient posterior inference on many observed datasets; see Figure 2 for a conceptual overview.
@@
2 Integrating amortized inference into the Bayesian workflow
Our adaptive workflow starts with neural network training to enable subsequent amortized inference on a large number of unseen datasets—typically well into tens of thousands. This training phase is conceptually identical to standalone amortized inference training [e.g., 62, 12]. For the inference phase, however, we develop a principled control flow that guides the analysis. Based on state-of-the-art diagnostics that are tailored to each step along the workflow, we propose decision criteria to select the appropriate inference algorithm for each observed dataset. In order to optimize the overall efficiency, our workflow contains mechanisms to reuse previous computations along the way.
@@ 2.1 Training phase: simulation-based optimization
In ABI, we train an amortized estimator $q_{\phi}$ which is parameterized by neural network weights $\phi$. This optimization objective is formalized as minimizing the Kullback–Leibler (KL) divergence between the true posterior $p(\theta\,|\,y)$ and the learned approximation $q_{\phi}(\theta\,|\,y)$ [59, 32, 62]:
$\phi=\arg\min_{\phi}\mathbb{E}_{p(y)}\left[\mathrm{KL}\left(p(\theta\,|\,y)\,\|\,q_{\phi}(\theta\,|\,y\right)\right]=\arg\min_{\phi}\mathbb{E}_{p(\theta)}\mathbb{E}_{p(y\,|\,\theta)}\left[-\log q_{\phi}(\theta\,|\,y)\right].$ (1)
Since most Bayesian models are generative by design, we can readily simulate $M$ synthetic tuples of parameters and corresponding observations from the joint probabilistic model,
@@
which results in the training set $\{(\theta^{(m)},y^{(m)})\}_{m=1}^{M}$ for optimizing Eq. missing1. Throughout this paper, we use coupling-based normalizing flows [18, 60] as a flexible conditional density estimator $q_{\phi}$. However, our proposed workflow is agnostic to the specific choice of neural network architecture used for amortization, as long as the model supports efficient sampling (see Section 2.2.1) and density evaluations (see Section 2.2.2).
@@
Since the neural network training algorithm hinges on simulated data, we cannot evaluate the amortized posterior approximator on real data just yet. However, we can easily simulate a synthetic test set $\{(\theta_{\star}^{(j)},y^{(j)})\}_{j=1}^{J}$ of size $J$ from the joint model via Eq. missing2. In this closed-world setting, we know which “true” parameter vector $\theta_{\star}^{(j)}$ generated each simulated test dataset $y^{(j)}$. A key diagnostic for evaluating the amortized posterior estimator is simulation-based calibration checking (SBC; [84, 69, 53]). Formally, SBC involves defining a test quantity $f:\Theta\times Y\rightarrow\mathbb{R}$ (e.g., marginal projections $\theta$ or the log likelihood $p(y\,|\,\theta)$), computing this statistic for the true data-generating parameter $\theta_{\star}^{(j)}$, and comparing it to the empirical distribution of the same statistic computed from amortized posterior draws given $y^{(j)}$ [53]. The rank of the true statistic within the posterior draws should be uniformly distributed if the amortized posterior estimator is well-calibrated. A complementary diagnostic to SBC is the parameter recovery check, where parameter estimates are compared against known ground-truth parameters [62, 64]. We refer to the appendix for details.
@@ 2.2 Inference phase: posterior approximation on observed datasets
Once the amortized estimator is capable of yielding sufficiently accurate posterior draws in closed-world settings (i.e., in-distribution), we use the pre-trained neural network to achieve rapid amortized posterior inference on a total of $K$ observed datasets $\{y_{\text{obs}}^{(k)}\}_{k=1}^{K}$. Recall that a given pre-trained amortized neural approximator may be perfectly suitable for some real datasets while it is utterly untrustworthy for others. Therefore, we want to assess on a per-dataset basis whether the amortized posterior draws are trustworthy and should be accepted, or whether we should proceed to a slower algorithm with stronger accuracy guarantees. The diagnostics in the inference phase are evaluated conditionally on each observed dataset, with the ultimate goal of determining whether the set of current posterior draws is acceptable for that specific dataset.
@@ 2.2.1 Step 1: Amortized posterior draws
We want to exploit the rapid sampling capabilities of the amortized posterior approximator $q_{\phi}$ as much as possible, as long as the sampled posteriors are reliable according to a set of principled diagnostics. Therefore, the natural first step for each observed dataset $y_{\text{obs}}^{(k)}$ is to query the amortized posterior and sample $S$ posterior draws $\hat{\theta}_{1}^{(k)},\ldots,\hat{\theta}_{S}^{(k)}\sim q_{\phi}(\theta\,|\,y^{(k)})$ in near-instant time (see Figure 2, first panel).
@@
Like other neural network based approaches, amortized inference may yield unfaithful results under distribution shifts [71, 91, 38]. To address this, we directly quantify whether an observed dataset $y_{\text{obs}}$ is atypical under the data-generating process $p(\theta,y)$. We first compute a low-dimensional summary statistic $s(y)\in\mathbb{R}^{d}$ for each dataset. [^1] The summary statistic can be either learned by the amortized estimator $q_{\phi}$ or be based on domain knowledge. The summary statistics from the training dataset $\{y^{(m)}\}_{m=1}^{M}$ are used to define a Mahalanobis distance. We compute this metric for each training dataset $y^{(m)}$ to establish a frequentist sampling distribution under the null hypothesis (i.e., of in-distribution datasets). Given a new dataset $y_{\text{obs}}$, we can now simply compare its empirical Mahalanobis distance to a predefined percentile threshold (e.g., $95^{\text{th}}$) of the sampling distribution (see Figure 3). In a nutshell, this is a sampling-based hypothesis test for distributional typicality, similar in spirit to the kernel-based test proposed by Schmitt et al. [71]. Since the amortized approximator has no guarantees nor known error bounds for data outside of the typical set of the joint model [19, 71, 23, 20], we propagate such atypical datasets to Step 2.
Figure 3: Illustration of our sampling-based hypothesis test that flags atypical datasets.
@@
In addition to the proposed OOD test, more sophisticated data-conditional diagnostics can further assess the accuracy of amortized posterior draws for individual datasets. Examples include posterior simulation-based calibration checking (posterior SBC; [83]) or the local classifier two-sample test (L-C2ST; [44]), to name a few. These diagnostics each offer distinct advantages and limitations; we refer to the appendix for a detailed discussion. By default, we recommend the OOD test for its simplicity, efficiency, and suitability as a first-line diagnostic.
@@ 2.2.2 Step 2: Pareto-smoothed importance sampling
In this step, we use a Pareto-smoothed sampling importance sampling (PSIS) scheme [88] to improve the quality of the amortized posterior draws of datasets which have previously been rejected (see Figure 2, second panel). Based on the amortized posterior draws from Step 1, PSIS computes importance weights $w_{s}^{\smash{(k)}}=p(y^{(k)}\,|\,\hat{\theta}_{s})\,p(\hat{\theta}_{s})/q_{\phi}(\hat{\theta}_{s}\,|\,y^{(k)})$ for each observed dataset $y^{(k)}$ (as in default importance sampling). Then, PSIS fits a generalized Pareto distribution, which in turn is used to smooth the tail of the weight distribution [88]. Finally, these smoothed importance weights are used for computing posterior expectations and for improving the posterior draws with the sampling importance resampling (SIR) scheme [68]. While the utility of standard importance sampling for improving neural posterior draws has previously been investigated [13], we specifically use the PSIS algorithm, which is self-diagnosing (see Diagnostics below) and therefore better suited for a principled workflow. Further details of PSIS are provided in the appendix.
@@
We use the Pareto-$\hat{k}$ diagnostic to gauge the fidelity of the PSIS-refined posterior draws. According to established guidelines [88, 97], Pareto-$\hat{k}\leq 0.7$ indicates good results, whereas $\hat{k}>0.7$ implies that the draws should be rejected and the respective datasets proceed to Step 3.
@@
The posterior estimator in ABI is typically mass-covering since it optimizes the forward KL divergence in Eq. missing1. When the neural network training is insufficient (e.g., small simulation budget or poorly optimized network), this may lead to overdispersed posteriors. Fortunately, this tends to err in the right direction, and PSIS can generally mitigate overdispersed mass-covering draws in low to moderate dimensions [15]. In contrast, variational inference typically optimizes the reverse KL divergence [67], which implies mode-seeking behavior that is less favorable for importance sampling.

### v3

@@ 1 Introduction
In many statistical modeling applications, from finance to biology and neuroscience, we often aim to infer unknown parameters $\theta$ from observables $y$ modeled as a joint distribution $p(\theta,y)$ (e.g., Raulo et al., 2023; Seaton et al., 2023; George et al., 2022; Landmeyer et al., 2020; Chen et al., 2019; Malén et al., 2022; Schneider et al., 2018; Tsilifis & Ghosh, 2022). The posterior $p(\theta\,|\,y)$ is the statistically optimal solution to this inverse problem, and there are different computational approaches to approximate this target distribution.
Markov chain Monte Carlo (MCMC) methods constitute the most popular family of posterior sampling algorithms and still remain the gold standard for modern Bayesian inference due to their theoretical guarantees and powerful diagnostics (Gelman et al., 2013; 2020). MCMC methods yield autocorrelated draws conditional on a fixed dataset $y_{\text{obs}}$. As a consequence, the probabilistic model has to be re-fit for each new dataset, which involves repeating the entire MCMC procedure from scratch. Modern implementations equip MCMC with state-of-the-art extensions, for example, through Hamiltonian dynamics (HMC; Neal, 2011), by minimizing the required tuning by users (NUTS; Hoffman & Gelman, 2014), or by parallelizing thousands of chains on GPU hardware (ChEES-HMC; Hoffman et al., 2021). The well-established Bayesian workflow (Gelman et al., 2020) leverages these tools in an iterative process of model specification, fitting, evaluation, and revision. While powerful, this approach becomes computationally burdensome when applied independently to large collections of datasets.
Differently, amortized Bayesian inference (ABI) aims to learn a direct mapping from observables $y$ to the corresponding posterior $p(\theta\,|\,y)$, using flexible function approximators such as deep neural networks (Cranmer et al., 2020; Radev et al., 2020; Greenberg et al., 2019; Papamakarios et al., 2021; Wildberger et al., 2023; Sharrock et al., 2024; Zammit-Mangion et al., 2025). Amortized inference typically follows a two-stage approach: (i) a training stage, where neural networks learn to distill information from the probabilistic model based on simulated examples of observations and parameters $(\theta,y)\sim p(\theta)\,p(y\,|\,\theta)$; and (ii) an inference stage where the neural networks approximate the posterior distribution for an unseen dataset $y_{\text{obs}}$ in near-instant time without repeating the training stage. In other words: The upfront training cost is amortized by negligible inference cost on arbitrary amounts of unseen test data. Owing to its reliance on simulated data, amortized inference in this form overlaps with simulation-based inference (Cranmer et al., 2020), which originated from posterior computations for models with intractable likelihood.
@@
However, amortized inference lacks the powerful diagnostics and gold-standard guarantees associated with MCMC samplers in the standard Bayesian workflow (Gelman et al., 2020). Yet, applying a standard workflow is computationally prohibitive at scale. In modern Bayesian computation, MCMC and ABI occupy different ends of a Pareto frontier (see Figure 1): the former provides reliable accuracy at high cost, while the latter offers near-instant inference speed with limited per-dataset reliability (Hermans et al., 2022; Schmitt et al., 2023; Lueckmann et al., 2021).
In this paper, we propose an adaptive workflow that consistently yields high-quality posterior draws while remaining computationally efficient. Our proposed workflow moves along the Pareto front, enabling fast-and-accurate inference when possible, and slow-but-guaranteed-accurate inference when necessary (see Figure 1). It combines the strengths of ABI and MCMC by incorporating diagnostic checks to guide inference decisions and reuse computations wherever possible. The resulting amortized Bayesian workflow therefore offers a principled, scalable, and diagnostic-driven approach for efficient posterior inference on many observed datasets; see Figure 2 for a conceptual overview. [^1] The software implementation is available at https://github.com/pipme/amortized-Bayesian-workflow. To summarize, our contributions are:
  • Design of—and systematic guidance through—an adaptive Bayesian workflow for accelerating Bayesian inference, which combines the strengths of amortized inference, importance sampling, and MCMC in a theoretically motivated and modular manner.
  • Empirical validation of the workflow and of its inference speedup, demonstrating the applicability of the workflow on both synthetic and large-scale, real-world problems.
2 Integrating amortized inference into the Bayesian workflow
Our adaptive workflow starts with neural network training to enable subsequent amortized inference on a large number of unseen datasets—typically well into tens of thousands. This training phase is conceptually identical to standalone amortized inference training (e.g., Radev et al., 2020; Cranmer et al., 2020). For the inference phase, however, we develop a principled control flow that guides the analysis. Based on state-of-the-art diagnostics that are tailored to each step along the workflow, we propose decision criteria to select the appropriate inference algorithm for each observed dataset. In order to optimize the overall efficiency, our workflow contains mechanisms to reuse previous computations along the way.
@@ 2.1 Training phase: simulation-based optimization
In ABI, a neural estimator $q_{\phi}$ with trainable parameters $\phi$ typically minimizes a strictly proper scoring rule $\mathcal{S}$ (Gneiting & Raftery, 2007; Pacchiardi & Dutta, 2022) in expectation over the joint model $p(\theta,y)=p(\theta)p(y\,|\,\theta)$,
$\phi=\operatorname*{arg\,min}_{\phi}\mathbb{E}_{(\theta,y)\,\sim\,p(\theta,y)}\left[\mathcal{S}\big(q_{\phi}(\cdot\,|\,y),\,\theta\big)\right].$ (1)
A popular choice is the logarithmic scoring rule, $\mathcal{S}\big(q_{\phi}(\cdot\,|\,y),\,\theta\big):=-\log q_{\phi}(\theta\,|\,y)$, which amounts to the forward Kullback-Leibler (KL) objective used for training normalizing flows in ABI (Greenberg et al., 2019; Radev et al., 2020). Score-based formulations that target a time-dependent gradient $\nabla_{\theta_{t}}\log p(\theta_{t}\,|\,y)$ are also possible (Sharrock et al., 2024; Gloeckler et al., 2024). Since most Bayesian models are generative by design, we can readily simulate $M$ synthetic tuples of parameters and corresponding observations from the joint probabilistic model,
@@
which results in the training set $\{(\theta^{(m)},y^{(m)})\}_{m=1}^{M}$ for optimizing Eq. 1. Throughout this paper, we use coupling-based normalizing flows (Durkan et al., 2019; Papamakarios et al., 2021) as a flexible conditional density estimator $q_{\phi}$ and the forward KL divergence as the training objective. However, our proposed workflow is agnostic to the specific choice of generative backbone used for amortization, as long as the model supports efficient sampling (see Section 2.2.1) and density evaluations (see Section 2.2.2).
@@
Since the neural network training algorithm hinges on simulated data, we cannot evaluate the amortized posterior estimator on real data just yet. However, we can easily simulate a synthetic test set $\{(\theta_{\star}^{(j)},y^{(j)})\}_{j=1}^{J}$ of size $J$ from the joint model via Eq. 2. In this closed-world setting, we know which “true” parameter vector $\theta_{\star}^{(j)}$ generated each simulated test dataset $y^{(j)}$. A key diagnostic for evaluating the amortized posterior estimator is simulation-based calibration checking (SBC; Talts et al., 2018; Säilynoja et al., 2022; Modrák et al., 2025; Yao & Domke, 2023). Formally, SBC involves (1) defining a test quantity $f:\Theta\times Y\rightarrow\mathbb{R}$ (e.g., marginal projections $\theta$ or the log likelihood $p(y\,|\,\theta)$), (2) computing this statistic for the true data-generating parameter $\theta_{\star}^{(j)}$, and (3) comparing it to the empirical distribution of the same statistic derived from amortized posterior draws given $y^{(j)}$ (Modrák et al., 2025). The rank of the true statistic within the posterior draws should be uniformly distributed if the amortized posterior estimator is well-calibrated.
We recommend assessing uniformity using the graphical approach by Säilynoja et al. (2022), which reveals the type of miscalibration present (e.g., bias or over-/under-dispersion) and is therefore useful for guiding improvements to amortized training. The choice of test quantity in SBC determines the sensitivity of the check; for example, the log-likelihood test quantity is typically more sensitive at detecting discrepancies than marginal projections (Modrák et al., 2025); using expressive neural classifiers is also possible (Yao & Domke, 2023). We further note a trade-off: imposing stricter criteria can improve the fidelity of the amortized estimator but will also tend to reject otherwise practically useful amortized estimators.
By default, we use marginal projections as the test quantities for SBC and complement SBC checking with parameter recovery checking, where parameter estimates are compared against known ground-truth parameters via direct visualization (Radev et al., 2020; 2023). Parameter recovery checking provides practical insight into whether the learned inverse mapping from $y$ to $\theta$ is effective and helps mitigate a known failure mode of SBC with marginal projections as test quantities, in which the posterior approximation simply recovers the prior. We refer to Appendix A for further details and corresponding pseudocodes.
Note.
Amortized inference lies at the intersection of Bayesian modeling and deep learning, unlocking massive potential for scalable posterior inference. However, this also comes with the practical challenges inherent to training deep neural networks. While a detailed treatment of neural architecture design and optimization exceeds the scope of this paper, practitioners can use established simulation-based inference libraries like sbi (Boelts et al., 2025) or BayesFlow (Radev et al., 2023), which provide modern plug-and-play components as well as sensible defaults for a wide range of applications. We summarize a set of best practices and actionable recommendations for training amortized posterior estimators in Appendix B.
@@ 2.2 Inference phase: posterior approximation on observed datasets
Once the amortized estimator is capable of yielding sufficiently accurate posterior draws in closed-world settings (i.e., in-distribution), we use the pre-trained neural network to achieve rapid amortized posterior inference on a total of $K$ observed datasets $\{y_{\text{obs}}^{(k)}\}_{k=1}^{K}$. Recall that a given pre-trained amortized neural estimator may be perfectly suitable for some real datasets while it is utterly untrustworthy for others. Therefore, we want to assess on a per-dataset basis whether the amortized posterior draws are trustworthy and should be accepted, or whether we should proceed to a slower algorithm with stronger accuracy guarantees. The diagnostics in the inference phase are evaluated conditionally on each observed dataset, with the ultimate goal of determining whether the set of current posterior draws is acceptable for that specific dataset.
@@ 2.2.1 Step 1: Amortized posterior draws
Figure 3: Illustration of our sampling-based hypothesis test that flags OOD datasets (to the right of the OOD cut-off).
We want to exploit the rapid sampling capabilities of the amortized posterior estimator $q_{\phi}$ as much as possible, as long as the sampled posteriors are trustworthy according to a set of principled diagnostics. Therefore, the natural first step for each observed dataset $y_{\text{obs}}^{(k)}$ is to query the amortized posterior and sample $S$ posterior draws $\hat{\theta}_{1}^{(k)},\ldots,\hat{\theta}_{S}^{(k)}\sim q_{\phi}(\theta\,|\,y^{(k)})$ in near-instant time (see Figure 2, first panel).
@@
Like other neural network approaches (Yang et al., 2024), amortized inference may yield unfaithful results under distribution shifts (Schmitt et al., 2023; Ward et al., 2022; Huang et al., 2023). To address this, we detect whether an observed dataset $y_{\text{obs}}$ is out-of-distribution (OOD) relative to the data-generating process $p(\theta,y)$. We first compute a low-dimensional summary statistic $s(y)\in\mathbb{R}^{d}$ for each dataset. [^2] The summary statistic can be either learned by the amortized estimator $q_{\phi}$ in the training phase or be based on domain knowledge. The summary statistics from the training dataset $\{y^{(m)}\}_{m=1}^{M}$ are used to approximate the Mahalanobis distance by estimating their empirical mean $\mu_{s}$ and covariance $\Sigma_{s}$. Then, for any test dataset $y$, its Mahalanobis distance to the training set is:
$D_{M}(y)=\sqrt{(s(y)-\mu_{s})^{\top}\Sigma_{s}^{-1}(s(y)-\mu_{s})}.$ (3)
We compute $\{D_{M}(y^{(m)})\}_{m=1}^{M}$ for all training datasets to establish a frequentist sampling distribution of distances under the null hypothesis (i.e., of in-distribution datasets). Given a new observed dataset $y_{\text{obs}}$, we compare its Mahalanobis distance $D_{M}(y_{\text{obs}})$ to the empirical distribution of training distances. We define the OOD rejection rule as:
$\text{Reject}_{\text{OOD}}(y_{\text{obs}})=\mathbb{I}\left\{D_{M}(y_{\text{obs}})>\text{Quantile}_{1-\alpha}\left(\{D_{M}(y^{(m)})\}_{m=1}^{M}\right)\right\},$ (4)
where $\alpha$ is by default set to 0.05 and we flag datasets whose Mahalanobis distances fall in the right $\alpha$ tail of the empirical training distances as out-of-distribution (see Figure 3). The type-I error rate $\alpha$ (false rejection) of this test can be set relatively high to obtain a conservative test that will flag many datasets for detailed investigation in further steps of our workflow.
In a nutshell, this is a sampling-based hypothesis test for distribution shifts, similar in spirit to the kernel-based test proposed by Schmitt et al. (2023). Since the amortized estimator has no guarantees nor known error bounds for data outside of the empirical support of the joint model $p(\theta,y)$(Elsemüller et al., 2024; Schmitt et al., 2023; Frazier et al., 2024; Elsemüller et al., 2025), we propagate such out-of-distribution datasets to Step 2. It is worth noting that a smaller Mahalanobis distance does not necessarily imply better posterior quality and that this OOD test is only intended to filter out datasets that are most likely to be problematic for the amortized estimator—specifically, those requiring extrapolation outside the ellipsoid defined by the training summary statistics.
@@
In addition to the proposed out-of-distribution test, more sophisticated data-conditional diagnostics can further assess the accuracy of amortized posterior draws for individual datasets and enhance the reliability of accepted amortized draws. Examples include posterior simulation-based calibration checking (posterior SBC; Säilynoja et al., 2025) or the local classifier two-sample test (L-C2ST; Linhart et al., 2023), to name a few. These diagnostics each offer distinct advantages and limitations, but typically require substantially more computation than the OOD test.
The choice to apply these additional diagnostics depends on context-specific factors, including the number of observed datasets, the relative computational cost of simulations versus likelihood evaluations, [^3] For example, if likelihood evaluations are relatively cheap, instead of applying sophisticated diagnostics in Step 1, it is often worthwhile to process directly to Step 2, where Pareto-smoothed importance sampling can offer more informative and powerful diagnostics. and the dimensionality of the observations. Ultimately, whether amortized posterior draws are deemed acceptable hinges on the accuracy requirements of the specific application. By default, we recommend the OOD test for its simplicity, efficiency, and suitability as a first-line diagnostic.
@@ 2.2.2 Step 2: Pareto-smoothed importance sampling
In this step, we use Pareto-smoothed importance sampling (PSIS) (Vehtari et al., 2024) to both improve and assess the quality of the amortized posterior draws of datasets which have previously been rejected (see Figure 2, second panel). Based on the amortized posterior draws from Step 1, PSIS computes importance weights $w_{s}^{\smash{(k)}}=p(y^{(k)}\,|\,\hat{\theta}_{s})\,p(\hat{\theta}_{s})/q_{\phi}(\hat{\theta}_{s}\,|\,y^{(k)})$ for each observed dataset $y^{(k)}$ (as in default importance sampling). Then, PSIS fits a generalized Pareto distribution to the largest importance weights, which in turn is used to smooth the tail of the weight distribution (Vehtari et al., 2024). Finally, these smoothed importance weights are used for computing posterior expectations and for improving the posterior draws with the sampling importance resampling (SIR) scheme (Rubin, 1988). While the utility of standard importance sampling for improving neural posterior draws has previously been investigated (Dax et al., 2023), we specifically use the PSIS algorithm, which is self-diagnosing (see Diagnostics below) and therefore better suited for a principled workflow. Further details of PSIS are provided in Appendix A.
@@
We use the Pareto-$\hat{k}$ diagnostic to gauge the fidelity of the PSIS-refined posterior draws. Pareto-$\hat{k}$ is the estimated shape parameter of the generalized Pareto distribution and quantifies the tail heaviness of the largest importance weights. According to Vehtari et al. (2024), for moderate sample size ($S>2000$), Pareto-$\hat{k}\leq 0.7$ indicates that PSIS estimates are reliable; [^4] For small sample size ($S<2000$), the threshold of Pareto-$\hat{k}$ is $\min(1-1/\log_{10}(S),0.7)$. when $\hat{k}>0.7$, the minimum sample size for obtaining a reliable Monte Carlo estimate through (Pareto-smoothed) importance sampling rapidly grows infeasibly large in practice, implying that the amortized posterior is a poor proposal for importance sampling correction and the corresponding dataset should be routed to Step 3. This $\hat{k}$ threshold is consistent with the established practice of using PSIS to improve and assess the quality of posterior approximations obtained from variational inference (Yao et al., 2018; Dhaka et al., 2021; Zhang et al., 2022).
@@
The posterior estimator in ABI is typically mode-covering since it optimizes the forward KL divergence in Eq. 1. When the neural network training is insufficient (e.g., small simulation budget or poorly optimized network), this may lead to overdispersed posteriors. Fortunately, this tends to err in the right direction, and PSIS can generally mitigate overdispersed mode-covering draws in low to moderate dimensions (Dhaka et al., 2021). In contrast, variational inference typically optimizes the reverse KL divergence (Rezende & Mohamed, 2015), which implies mode-seeking behavior that is less favorable for importance sampling.